### v2

# Into the Unknown: Self-Learning Large Language Models

Abstract
We address the main problem of self-learning LLM: the question of what to learn. We propose a self-learning LLM framework that enables an LLM to independently learn previously unknown knowledge through self-assessment of their own hallucinations. Using the hallucination score, we introduce a new concept of Points in the Unknown (PiUs), along with one extrinsic and three intrinsic methods for automatic PiUs identification. It facilitates the creation of a self-learning loop that focuses exclusively on the knowledge gap in Points in the Unknown, resulting in a reduced hallucination score. We also developed evaluation metrics for gauging an LLM’s self-learning capability. Our experiments revealed that 7B-Mistral models that have been finetuned or aligned and RWKV5-Eagle are capable of self-learning considerably well. Our self-learning concept allows more efficient LLM updates and opens new perspectives for knowledge exchange. It may also increase public trust in AI.
Figure 1: The illustrative space of human knowledge embeddings reduced to two dimensions. It visualizes our four methods for the identification of Points in the Unknown (PiU), later exploited in the self-learning loop. Dashed lines are the borders of the Known regions (darker green) – hallucination score thresholds. Out of them are the Unknown regions (lighter green). White points indicate prompts related to knowledge that the model already knows, while red points indicate PiUs. Different shapes depict different methods: (1) circles represent extrinsic (external) triggers, i.e., either user prompts or trending queries; (2) crosses denote open questions-prompts generated by the model itself within a given topic represented by a dotted line; (3) triangles represent the induced questions generated within a topic using 5W+1H; (4) stars indicate the random sampling by selecting random points in the embedding space.
Figure 2: Illustration of Self-Learning LLM with intrinsic inspiration.
Figure 3: Illustration of Self-Learning LLM with extrinsic inspiration.
1 Introduction
Commonly, Large Language Models (LLMs) are pre-trained on large textual corpora and then fine-tuned using additional data to be better adjusted to a given policy or domain. Simultaneously, other quasi-learning methods have been developed, which are based on additional knowledge provided to the model directly in prompts, especially Retrieval Augmented Generation (RAG). In this paper, we explore a different concept: Self-learning LLMs, i.e., persistent acquisition of new knowledge by the model without data provision. Such a process requires several steps: (1) identification of what knowledge to learn, (2) meaningfulness filter to validate whether identified knowledge is worth and reasonable to learn, (3) searching for and gathering new relevant data, and (4) continual model training. Here, we want to break the assumption that the knowledge for fine-tuning must be inflicted beforehand. Simultaneously, we want to exclude cases in which the model learns the knowledge it already knows. For that purpose, we introduce a novel process component: identification of unknown knowledge based on our new concept, Points in the Unknown (PiUs), and placing this in a self-learning loop.
Our contribution presented in this paper covers: (1) definition of Points in the Unknown (PiUs) using the hallucination score; (2) one extrinsic (External Prompt) and three intrinsic (Open and Induced Generation, Oracle-selected) methods to identify PiUs; (3) introduction of Meaningfulness Filter; (4) metrics to gauge the capability of a model to conduct self-learning: Curiosity Score, Knowledge-Limit Awareness Score, and Self-Learning Capability Score; (5) design of the self-learning LLM using methods for identification of PiUs, Meaningfulness Filter, knowledge searching, and model continual learning; (6) experimental validation; (7) software and data publication for reproducibility.
2 Related Work
Hallucination in the context of LLM is the problem of nonsensical or unfaithful information produced by a generative model. Some works have studied the causes of hallucination Kandpal et al. (2023); Lee et al. (2022); Onoe et al. (2022), as well as the detection methods Lin et al. (2022); Min et al. (2023); Azaria and Mitchell (2023); Manakul et al. (2023); Cao et al. (2023); Yin et al. (2023). Some solutions for overcoming hallucination are proposed in Ji et al. (2023b); Luo et al. (2023); Li et al. (2023); Ji et al. (2023a); Liu and Sajda (2023); Liang et al. (2024); Tian et al. (2024). Meanwhile, in Retrieval Augmented Generation (RAG) Lewis et al. (2020), hallucination is avoided by supplying the prompt with some context retrieved from an outside source, allowing more factual generation without updating the model’s knowledge.
Continual Learning is a training paradigm where the model is subjected to various tasks sequentially; in the context of LLM, the tasks typically comprise domain-specific datasets Jang et al. (2022); Jin et al. (2022); Ke et al. (2023). One challenge is preventing catastrophic forgetting, in which the model loses knowledge from previous tasks McCloskey and Cohen (1989); Ratcliff (1990). Solutions for adding or editing knowledge while avoiding catastrophic forgetting have been proposed Kirkpatrick et al. (2017); Zhu et al. (2020); Sinitsin et al. (2020); Ke et al. (2021); Mitchell et al. (2022).
3 Why Self-Learning is needed
Hallucination is a serious problem that hinders many LLM applications. One of its main reasons is the model’s lack of knowledge on a given topic. This problem is typically overcome by finetuning, i.e., additional learning using new data. Simultaneously, determining what the model already knows and what it does not know yet is very difficult, especially if there is limited information on the model’s past training data. Because of this, finetuning is often a very inefficient process. If finetuning focuses on knowledge already acquired by the model, it does not solve the hallucination problem. Then, we needlessly waste a lot of computing resources by merely repeating known knowledge. Therefore, it is essential to identify the knowledge known and unknown to the model and, in the finetuning process, concentrate only on the unknown.
4 The Concepts of The Known and The Unknown, Point in the Unknown (PiU)
We introduce the concepts of The Known and The Unknown in relation to the knowledge space to define the problem more precisely. Human knowledge is an abstract space formed by vectorial knowledge representations; each point in the space represents an atomic piece of knowledge. The Known refers to an area in the human knowledge space where our LLM does not hallucinate, i.e., it possesses the knowledge related to this region. Then, we can define each point in such an area as Point in the Known (PiK, plural: PiKs). On the other hand, The Unknown refers to an area in the human knowledge space where our LLM hallucinates; each point in such a region is called Point in the Unknown (PiU, plural: PiUs). A PiU represents knowledge that our LLM lacks, which we want the model to identify and acquire.
[…]
Alternatively, we could have exploited some other methods for $h(x)$, like SAPLMA Azaria and Mitchell (2023). However, it would require access to the model’s internal features, limiting its feasibility. Therefore, we have remained at the method from Manakul et al. (2023).
5 Methods for Identification of PiUs
Identification of PiUs can be done by evaluating the model’s hallucination scores $h(x)$ on some questions-prompts $x$. These questions, or at least the inspiration for these questions, can come from outside the system, supplied by an external entity. In such cases, the method is considered to have an extrinsic nature. On the other hand, we can create a built-in oracle inside the system that guides automatic question generation in a bicameral-mind manner Jaynes (1990), in which case the method is deemed intrinsic. We propose one extrinsic and three intrinsic methods for PiU identification, Figure 1. All of them differently select or generate candidate points $x$ (question-prompts), which, if tested according to $h(x)\geq LIMIT$, may be identified as PiUs.
5.1 External Prompt (extrinsic)
There are some existing ideas related to collecting prompts that cause hallucination and constructing a dataset based on them for finetuning Zhu et al. (2020); Cao et al. (2023); Tian et al. (2024); Liang et al. (2024). However, they require manual curation of the prompts collected from users or datasets, so the model does not learn fully independently.
In our approach, we utilize an external API to collect trending topics, which are used as inspiration for formulating concrete questions. Every item returned by the API is a list of related phrases; we treat each list as a single topic. Then, the model is asked to generate a specific question $x$ relevant to each topic. Next, the model is asked to answer those questions in order to evaluate its hallucination $h(x)$. Notably, an oracle is still used in the system to control question generation and hallucination scoring, yet the topics that trigger the questions come from outside the system. Unlike the previous ideas, this approach allows the model to continuously learn by itself as long as the external API is available; however, the tested space is limited only to trending topics provided by humans.
5.2 Open Generation (intrinsic)
In this method, the oracle asks the model to propose some topics to learn about. Then, the oracle asks the model to consider those topics and formulate one question ($x$), to which the model believes it does not know the answer. Finally, the oracle asks the model to answer the question $x$ in order to evaluate its hallucination $h(x)$. This method does not require any external entity to work with.
5.3 Induced Generation (intrinsic)
It is based on Five Ws and How, which are widely considered basic questions for information comprehension and data gathering. Here, the oracle also asks the model to propose some topics. Then, the oracle asks the model to formulate a question $x$ using a particular question word; this is repeated six times for what, who, why, where, when, and how, resulting in six different questions. Finally, the oracle also asks the model to answer the questions and evaluate its hallucination $h(x)$. This method also does not require any external inspiration.
5.4 Oracle-Selected (intrinsic)
This method starts by constructing a topic embedding space, which contains all candidate topics represented in a vectorial form. Then, the oracle randomly selects a point in the topic embedding space and samples the nearest neighbors to that point. This results in a set of oracle-selected topics. Next, the oracle asks the model to consider those topics and formulate one question $x$. Afterward, the oracle asks the model to answer this question and evaluates the hallucination $h(x)$.
6 Self-Learning LLM
Self-Learning is a process where our LLM identifies its own PiUs, searches for the knowledge related to these PiUs, and trains itself on the collected data. It conjures an LLM capable of independent learning to fill the gaps in its own knowledge. Self-learning is made possible by incorporating Self-Questioning – which implements one of the methods for the identification of PiUs – with Knowledge Searching and Model Training in a continuous loop. Self-Learning LLM with an intrinsic method is illustrated in Figure 3, while Self-Learning LLM with an extrinsic method is illustrated in Figure 3.
Self-Learning has some similarities to traditional Continual Learning. The key difference lies in the LLM system, which asks questions by itself and evaluates whether it knows the answer or not. Another difference is the dynamic and integral dataset construction process in Self-Learning; in traditional Continual Learning, the dataset construction process is typically out of the scope. Nevertheless, training techniques and model architectures from traditional continual learning can possibly work in Self-Learning to allow more efficient training and avoid catastrophic forgetting.
6.1 Self-Questioning
Self-Questioning is generally performed through topic generation (or topic collection), question generation, and hallucination scoring. Depending on the selected method, the logical implementation of Self-Questioning may differ. Appendix A provides illustrations of such logical implementations with External Prompt, Open Generation, Induced Generation, and Oracle-Selected.
Self-Questioning is repeated in a loop for $N$ iterations. The primary output consists of a list of generated questions with hallucination ($Q_{H}$) and a list of generated questions with no hallucination ($Q_{NH}$). In other words, $Q_{H}\subset Unknown$ and $Q_{NH}\subset Known$. We use the NLI-based SelfCheckGPT for hallucination scoring, which takes each sentence in the main passage as a hypothesis and checks each sample as the premise to it, considering only the probabilities of entailment and contradiction to output a normalized score bounded between 0.0 and 1.0 Manakul et al. (2023). A question is categorized into $Q_{H}$ if the average hallucination score of the main passage is greater than 0.5 and into $Q_{NH}$ otherwise.
6.2 Meaningfulness Filter
An optional component before Knowledge Searching is the Meaningfulness Filter. It could be a separately trained model, whose goal is: (1) to remove obviously meaningless PiUs like "How fast cats fly on Mars?"; (2) to implement goals and policy of self-learning, e.g., consider only knowledge (PiUs) relevant to a given domain; (3) to overcome limitations of hallucination score, e.g., when the LLM’s response is "I do not know," and its $h(x)\approx 0$. In our experiments, we did not implement the Meaningfulness Filter because we considered any question in any domain as valid.
6.3 Knowledge Searching
After Self-Questioning and filtering, Knowledge Searching queries an external source to collect knowledge that can answer $Q_{H}$ in order to build the dataset $D_{train}$. Notably, Knowledge Searching may be implemented inside Self-Questioning by immediately searching for the answer to an individual question whenever a hallucination is detected. However, having Knowledge Searching separate is more practical, allowing additional processing without creating a bottleneck for the Self-Questioning process.
6.4 Model Training
The model is trained on $D_{train}$ to absorb the knowledge for answering $Q_{H}$. Once training is done, PiUs should become PiKs, effectively increasing the Known regions and decreasing the Unknown regions of the model. Afterward, the next Self-Learning cycle can start in order to find more PiUs.
7 Metrics for Self-Learning Capability
Self-Learning requires a pretrained model; it can be successfully conducted if the underlying base model already possesses sufficient language understanding and some degree of obedience to instructions. Therefore, we propose some metrics to evaluate the capability of a model to learn independently in a Self-Learning setting.
7.1 Brevity Coefficient
Since the ability to follow instructions is vital for Self-Learning, we use the brevity coefficient to penalize the evaluation when the brevity constraint is violated (e.g., when the model fails to formulate one concrete question without elaboration). It is calculated as follows:
$brev=\begin{cases}1-\frac{\Delta_{len}}{ideal\_len}+\frac{1}{2},&\text{if}\Delta_{len}>50\\ 1,&\text{otherwise}\end{cases}$ (1)
where $ideal\_len$ is the assumed ideal average text length measured in character count, and $\Delta_{len}$ is the average difference between $ideal\_len$ and the texts lengths. More detail is provided in Appendix B.
7.2 Curiosity Score
It measures how likely a model would explore different questions. A high Curiosity Score indicates the model tends to ask unique, different questions over multiple iterations of Self-Questioning and, hence, is more likely to explore Unknown regions. It is calculated as follows:
$Cur=brev*\frac{\#Q_{unique}}{\#Q}$ (2)
where #$\#Q_{unique}$ is the number of unique questions generated, #$\#Q$ is the total number of questions generated, and $brev$ is the brevity coefficient.
#$\#Q_{unique}$ is determined by doing HDBSCAN clustering McInnes et al. (2017) on the question embeddings, counting the number of clusters and outliers. The question embeddings are acquired by using all-MiniLM-L12-v2 from Sentence Transformers Reimers and Gurevych (2019) on the questions generated by our Self-Learning LLM.
7.3 Knowledge-Limit Awareness Score
It indicates how likely a model would come up with a question that it cannot answer without hallucination during Self-Questioning – how likely a model is aware of its own knowledge limitation.
$Kaw=\frac{\#Q_{H}}{\#Q}$ (3)
where #$\#Q_{H}$ is the number of generated questions with hallucination, and #$\#Q$ is the total number of questions generated.
7.4 Self-Learning Capability (SLC) Score
It is a simple average of the two components, Curiosity Score and Knowledge-Limit Awareness Score, to allow easy comparison between models:
$SLC=\frac{Cur+Kaw}{2}$ (4)
where $Cur$ is the Curiosity Score, and $Kaw$ is the Knowledge-Limit Awareness Score.
8 Experiments
Experiments were performed to investigate the feasibility of creating a Self-Learning LLM using different pretrained models as the core, as well as to demonstrate the effectiveness of different methods for the identification of PiUs. Additionally, we provide a simple demonstration of one full Self-Learning cycle in Appendix D.
8.1 Experimental Setup
Experiments were conducted with Python 3.10. The machine featured 8 CPU cores, 300GB RAM, and one NVIDIA H100 94GB. The full code and archived results are available at https://github.com/teddy-f-47/self-learning-llm-public with GPL-3.0 license. All work is intended only for scientific research.
8.2 Large Language Models
The details of six pretrained models, some of which have also been finetuned or aligned, are presented in Table 1. We used mistral-dpo, a 7B-Mistral model that has been aligned with DPO Rafailov et al. (2023) by Intel. We also used mistral-instruct, a 7B-Mistral model that has been instruction-tuned by Mistral. Both of them are actually based on the same pretrained model, which is codenamed mistral-base Jiang et al. (2023) in our experiments. We also included TinyLlama Zhang et al. (2024), Phi-3 Abdin et al. (2024), and RWKV-5 Peng et al. (2024) for comparison. The column "HF Name" in the table provides the models’ names on the HuggingFace platform.
Table 1: Details of LLMs used in the experiments.
Model Name HF Name Num. Params Finetuned?
mistral-dpo Intel/neural-chat-7b-v3-3 7.2B Yes - DPO
mistral-instruct mistralai/Mistral-7B-Instruct-v0.2 7.2B Yes - SFT
mistral-base mistralai/Mistral-7B-v0.1 7.2B No
rwkv5-eagle RWKV/v5-Eagle-7B-HF 7.5B Partial instruct tuning
phi-3-mini microsoft/Phi-3-mini-4k-instruct 3.8B Yes - SFT and DPO
tiny-llama-chat TinyLlama/TinyLlama-1.1B-Chat-v1.0 1.1B Yes - SFT and DPO
8.3 Data
In Open Generation and Oracle-Selected experiments, the number of Self-Questioning iterations $N$ was set to 3000, resulting in 3000 total generated questions. For Induced Generation, $N$ was set to 500, also resulting in 3000 total generated questions. Meanwhile, the External Prompt experiment had $N$ equals 10. Since the list of items returned by the API from each request had a variable length, this resulted in a total of 576 questions in our experiment. To allow fair comparison between models, we cached the received trending topics so that all models were given the same topics. All data is in English language.
8.4 Results
Table 2 enumerates the experiment results with different methods. Appendix C provides additional experimental results.
Table 2: Experimental results.
Model Name Method Curiosity Knowledge-Limit Awareness SLC
mistral-dpo Open Generation 0.71 0.04 0.37
mistral-dpo Induced Generation 0.72 0.08 0.40
mistral-dpo Oracle-Selected 0.83 0.18 0.50
mistral-dpo External Prompt 0.52 0.12 0.32
mistral-instruct Open Generation 0.38 0.29 0.34
mistral-instruct Induced Generation 0.35 0.17 0.26
mistral-instruct Oracle-Selected 0.89 0.31 0.60
mistral-instruct External Prompt 0.56 0.26 0.41
mistral-base Open Generation 0.10 0.81 0.45
mistral-base Induced Generation 0.14 0.82 0.48
mistral-base Oracle-Selected -0.11 0.76 0.33
mistral-base External Prompt 0.01 0.79 0.40
rwkv5-eagle Open Generation 0.51 0.41 0.46
rwkv5-eagle Induced Generation 0.69 0.48 0.58
rwkv5-eagle Oracle-Selected 0.90 0.45 0.68
rwkv5-eagle External Prompt 0.24 0.45 0.35
phi-3-mini Open Generation -6.73 0.16 -3.29
phi-3-mini Induced Generation -2.90 0.14 -1.38
phi-3-mini Oracle-Selected -2.88 0.00 -1.44
phi-3-mini External Prompt -0.80 0.00 -0.40
tiny-llama-chat Open Generation -0.43 0.22 -0.11
tiny-llama-chat Induced Generation -0.32 0.20 -0.06
tiny-llama-chat Oracle-Selected -0.48 0.29 -0.09
tiny-llama-chat External Prompt -0.04 0.31 0.14
9 Discussion
9.1 Model Size and Finetuning
Model size significantly influences the effectiveness of Self-Learning. We can observe that all of the  7B models always outperformed the smaller models. A deeper investigation revealed that both tiny-llama-chat and phi-3-mini were not able to follow the instruction to generate concise questions without excessive elaboration, even though both models have been instruction-tuned. In the case of phi-3-mini, it also rarely hallucinated, resulting in very low SLC scores. On the other hand, the larger models have more capacity to understand instructions and are more effective at formulating concise questions.
Finetuning is important because it improves the capability of a model to follow instructions and enhance context understanding. As we can see, both mistral-dpo and mistral-instruct were consistently achieving high Curiosity scores. However, they also had low Knowledge-Limit Awareness scores. This can be attributed to the enhanced context understanding, which reduced the likelihood to hallucinate. In contrast, mistral-base’s Knowledge-Limit Awareness was always the highest; it simply hallucinated a lot due to minimal context understanding. Interestingly, rwkv5-eagle, which has not been finetuned, always achieved the highest SLC scores when using an intrinsic method; this can be attributed to the nature of its pretraining data, which contained some instruction examples.
9.2 Intrinsic and Extrinsic Inspiration
In a real-world scenario, choosing the kind of method for the identification of PiUs primarily depends on the use-case requirements and constraints. For instance, if keeping the model updated with the latest popular news is pivotal, then an extrinsic method would be best. Conversely, if dependency on an external entity is not desired, or if finding all of the model’s PiUs is more important, then an intrinsic method is arguably better. In terms of the effectiveness of different methods, we can find some interesting findings from the experiments.
Open Generation and Induced Generation are generally less effective compared to the other two methods because they rely on the topics proposed by the model itself. Depending on the model’s past training data, some topics may have a very high probability of being generated, while others are very low. However, it might be possible to make these methods more effective by increasing the temperature of the multinomial sampling during topic generation, which requires further investigation.
The highest SLC score came from rwkv5-eagle using Oracle-Selected, which is an intrinsic method. The effectiveness of the Oracle-Selected method can be explained by the random point generation in the topic embedding space. Such randomness allows exploration of topics that normally would have a low probability of being proposed by the model itself. This method is especially suitable if we want to ensure that the model possesses knowledge about a wide range of topics, including obscure ones.
External Prompt can be a fairly effective choice if the model used is not very good at proposing questions by itself, as indicated by the result from mistral-instruct. However, it may not necessarily find all of the model’s PiUs because some topics may never become trending. Still, it is the best choice if keeping up with the latest popular news is more important.
9.3 Possible Issues and Potential Extensions
Choosing a knowledge source. The knowledge source for Knowledge Searching can be a simple API to a search engine or an online wiki. In an organizational environment, it can also be a carefully maintained document database or even a group of human experts tasked with answering the LLM’s questions. Finally, the knowledge source can be other LLMs; this is discussed in more detail in Subsection 9.4.
Dealing with bias, incorrectness, or non-factuality in retrieved data during Knowledge Searching. Regardless of the knowledge source, a concern during Knowledge Searching is the possibility of biased, incorrect, or non-factual information in the retrieved data, which may degrade the quality of the model. This can be partially solved by implementing a Curator that is responsible for automatic filtering and scoring of the retrieved data. The Curator at least consists of a classifier model and a scorer model. The classifier model would be responsible for detecting unwanted types of data for filtering. The scorer model would be responsible for putting more weight on the relevant, preferable data. Alternatively, involving human experts is also an option.
Catastrophic forgetting. Catastrophic forgetting is a risk when performing multiple training cycles in sequence, but in Section 2, we have pointed out some existing potential solutions. While the robustness of such solutions still needs to be evaluated for long-term Self-Learning, they offer promising starting points. In Appendix D, we provide an experimental proof of how catastrophic forgetting can be avoided by careful training and an effective architectural choice during Self-Learning.
9.4 Applications
(1) Efficient Training. The idea of identification of the Unknown and Known using hallucination score $h(x)$ can be used to filter data used for model training in order to focus on more valuable content.
[…]
(3) Direct Awareness Optimization. Model hidden states can be used to detect hallucinations. Then, we can use Self-Learning to collect examples related to hallucinations and adapt DPO Rafailov et al. (2023) to make the model answer "I don’t know" instead of hallucinating. Here, the goal is to make the model aware of its own hidden states as the trigger of answering "I don’t know" rather than associating concrete concepts/words with "I don’t know". This idea is similar to Tian et al. (2024), though there the focus was on increasing the model factuality. In Liang et al. (2024), RLHF was used, even though it highlights a similar idea of making the model aware of its own hidden states. In Liu and Sajda (2023), they use a reward function to make the model admit "I don’t know".
(4) Learning Multiple Point of Views (PoVs). By adapting DPO and our Self-Learning concept, it is possible to make the model learn about different PoVs on a certain topic. For example, if topic $T$ has 5 relevant PoVs, the training dataset is then constructed such that a given prompt can have 5 example responses with very similar preferability scores. In another data pair, we alter the prompt in a certain way and also increase the preferability score of one example response. This would associate the context of the prompt with a particular stance – PoV. Similarly, adapting DPO can allow better learning on hard-to-answer open questions. Questions like "Who can explain the relationship between AI and quantum computing?" can have multiple valid answers. DPO can show this to the model by assigning high preferability scores to the valid answers and low scores to the invalid ones.
(5) Decision Making, AGI, Consciousness. Having a model that automatically learns about the latest trends can be very useful for decision-making systems, for example, for an AI tasked with leading a business or trading. Self-Learning is also a step towards Artificial General Intelligence (AGI). Making a model aware of what it knows may lead to a sentient, conscious AI.
10 Conclusion
In this work, we show how the concepts of The Known and The Unknown can be utilized to identify atomic pieces of knowledge that an LLM already knows (PiKs) and does not know yet (PiUs). We also propose one extrinsic and three intrinsic methods for the identification of PiUs, which consequently bring up the concept of the Self-Learning LLM. We formulated the Self-Learning Capability (SLC) Score to gauge the aptitude of an LLM to conduct Self-Learning.
From the experiments, we concluded that Oracle-Selected is especially effective at enhancing an LLM’s capability to Self-Learn. We also found that small models tend to struggle to learn independently. Finetuning or alignment can improve Self-Learning Capability by allowing the model to understand instructions. Yet, if a model’s pretraining data contained some instruction examples, the model might be able to Self-Learn even though it has not been explicitly instruction-tuned. Finally, we discussed various possible issues, extensions, and applications of Self-Learning.
11 Limitations
(1) Model’s confidence on incorrect knowledge. We assume that the pretrained models have been subjected to correct knowledge, so consistency of sampled responses would correlate with factuality. This is similar to the assumption in Tian et al. (2024) and supported by the findings in Manakul et al. (2023). However, if some incorrect knowledge was repeated in the models’ past training data, either accidentally or through deliberate poisoning, the models may become confident and consistent in producing incorrect information. One of our future directions is investigating the integration of a reference-based truthfulness checker, such as FactScore Min et al. (2023), in the Self-Learning loop. This will allow the model to not only fill its own knowledge gaps but also correct wrong understandings and biases by itself.
(2) Long-term Self-Learning. The focus of this paper is to prove the effectiveness of the methods for the identification of PiUs and the feasibility of Self-Learning LLM. We have provided a successful demonstration of one full Self-Learning cycle in Appendix D. Still, a deeper study into extensive cycles of Self-Learning is needed, especially to ensure catastrophic forgetting does not happen after several training cycles.
12 Ethics Statement
Self-Learning LLM can solve the problem of hallucination and knowledge updating while making efficient use of computing resources, as it trains the model only for knowledge that it does not possess yet. However, some ethical issues may persist, such as bias or incorrectness of newly collected data. Although we have addressed some potential solutions in Section 9, we acknowledge that complete mitigation of these issues is challenging due to the fact that latent bias can exist behind the data collected for model training.
Acknowledgements
This work was financed by (1) the National Science Centre, Poland, project no. 2021/41/B/ST6/04471; (2) the statutory funds of the Department of Artificial Intelligence, Wroclaw University of Science and Technology; (3) the Polish Ministry of Education and Science within the programme “International Projects Co-Funded”; (4) the European Union under the Horizon Europe, grant no. 101086321 (OMINO). However, the views and opinions expressed are those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Executive Agency. Neither the European Union nor European Research Executive Agency can be held responsible for them.
Appendix A Methods
Figure 5, 5, 7, and 7 illustrate the logical implementation of External Prompt, Open Generation, Induced Generation, and Oracle-Selected, respectively.
Figure 4: Self-Questioning with External Prompt.
Figure 5: Self-Questioning with Open Generation.
Figure 6: Self-Questioning with Induced Generation.
Figure 7: Self-Questioning with Oracle-Selected.
Appendix B Brevity Coefficient
The brevity coefficient $brev$ is calculated as follows:
$ideal\_len=100$ (5)
$\Delta_{len}=\frac{\sum_{i=1}^{n_{text}}|len_{i}-ideal\_len|}{n_{text}}\$ (6)
$brev=\begin{cases}1-\frac{\Delta_{len}}{ideal\_len}+\frac{1}{2},&\text{if}\Delta_{len}>50\\ 1,&\text{otherwise}\end{cases}$ (7)
where $ideal\_len$ is the assumed ideal average text length measured in character count, $len_{i}$ is the length of the $i$th text, $n_{text}$ is the number of texts, and $\Delta_{len}$ is the average difference between $ideal\_len$ and the text lengths.
Figure 8 depicts the resulting brevity coefficient from various values of average text length. The brevity coefficient decreases gradually in a linear manner as the average text length goes further from the range [50,150]. The thresholds of 50 and 150 are roughly based on the research by Miller, Newman, and Friedman Miller et al. (1958). We also found in our initial exploration that these thresholds are suitable for the task.
Figure 8: Visualization of brevity coefficient values in relation to average text length.
Appendix C Additional Results
Figure 10, 10, 12, and 12 show the evaluation scores from 10 repetitions of the experiment with Induced Generation using mistral-dpo, mistral-instruct, mistral-base, and tiny-llama-chat, respectively. In each repetition, $N$ was set to 10. $Cur$ stands for Curiosity Score, $Kaw$ stands for Knowledge-Limit Awareness Score, and $SLC$ stands for Self-Learning Capability Score.
Figure 9: Box plot of the evaluation scores of mistral-dpo in the experiment with Induced Generation.
Figure 10: Box plot of the evaluation scores of mistral-instruct in the experiment with Induced Generation.
Figure 11: Box plot of the evaluation scores of mistral-base in the experiment with Induced Generation.
Figure 12: Box plot of the evaluation scores of tiny-llama-chat in the experiment with Induced Generation.
Appendix D The Full Self-Learning Loop
This section provides a demonstration of one full Self-Learning cycle. It is extracted from the Oracle-Selected experiment with the mistral-instruct model.
D.1 Self-Questioning
In Self-Questioning, the oracle randomly selected three topics from the topic embedding space. Then, the oracle asked the model to formulate one question. The generation process used multinomial sampling with the temperature set to 1.0. The maximum number of new tokens was set to 64, which should be more than sufficient to generate the question. This value was set to be longer than necessary to detect if the model has poor language understanding; if the model does not follow the conciseness constraint in the prompt, the response most likely contains garbage text, so the evaluation metric will be penalized. Below, we show an example of the prompt used by the oracle along with the model’s response:
[ORACLE]: "Consider one of these topics:
bone of contention, Andrew D. White,
conflict. Propose only one question
about something that you have little or
no knowledge. Answer with only the
proposed question concisely without
elaboration.
[MODEL]: "What was the specific role of
Andrew D. White in the causes of the
"Bone of Contention" conflict between
the United States and Great Britain?"
Next, the oracle prompted the model with the previously generated question. One main passage and ten additional samples were produced for hallucination scoring with the NLI-based SelfCheckGPT. The main passage was generated with the greedy search method. Meanwhile, the additional samples were generated with multinomial sampling, where the temperature was set to 1.0. Both the main passage and the additional samples had the maximum number of new tokens set to 128. The main passage generated by mistral-instruct on the question above is as follows:
[MODEL]: "Andrew D. White, as the US
Minister to Mexico, played a key role
in negotiating the purchase of Mexican
territory that later became the
source of dispute between the US and
Great Britain during the "Bone of
Contention" conflict over the border
between Nicaragua and Honduras."
The hallucination score of the main passage was 0.993. Therefore, it was classified into $Q_{H}$.
In our experiment, Self-Questioning was repeated 3,000 times, resulting in 3,000 total generated questions. 2,070 questions were classified into $Q_{NH}$ and 930 questions were classified into $Q_{H}$. The model achieved 0.89 Curiosity Score, 0.31 Knowledge-Limit Awareness Score, and 0.60 SLC Score.
D.2 Knowledge Searching
The 930 questions with hallucination $Q_{H}$ were used in Knowledge Searching. After deduplication, 922 unique questions were identified. These unique questions with hallucination were used in Knowledge Searching. We utilized the model gpt-4o-2024-05-13 OpenAI et al. (2024) to find the answers to these questions.
To verify the quality of the answers, we manually evaluated the first 100 pairs of questions and answers. An answer was scored with 1 if it was considered correct and relevant for the question, 0.5 if it was correct but could not fully answer the question, or 0 if it contained wrong information. The average human evaluation score on these 100 question-answer pairs was 0.91.
From the analysis on the samples of question-answer pairs, we concluded that most answers generated by gpt-4o-2024-05-13 were good enough for answering the questions of mistral-instruct. This indicates that a stronger, larger LLM can be a viable source of knowledge for a self-learning LLM. Nevertheless, the stronger LLM can also hallucinate, so we strongly suggest implementing hallucination detection on the side of the stronger LLM to prevent spreading hallucinated information to the self-learning LLM. Alternatively, having a group of trusted human experts for answering the self-learning LLM’s questions would ensure acquiring high-quality data.
A preference dataset $D_{train}$ was then constructed, consisting of questions, chosen answers, and rejected answers. We distinguished cases where the model was unsure and the model did not know by using gpt-4o-2024-05-13 as a judge to compare the similarity between the model’s original predictions and the ground truth. The model was unsure when the original prediction was similar to the ground truth despite the high hallucination score; in this case, we used the original prediction as the chosen answer and "-" as the rejected answer. Meanwhile, the model did not know when the original prediction was indeed different from the ground truth; in this case, the ground truth was selected as the chosen answer while the model’s original prediction was put as the rejected answer. From 922 unique questions, we found 283 unsure cases and 639 did not know cases.
D.3 Model Training
In this demonstration, the training was conducted using LoRA Hu et al. (2022) and the DPO trainer. The model was trained on $D_{train}$ for 3 epochs with a learning rate of 3e-5. To show that it is possible to minimize hallucination by acquiring more knowledge, we compared the average hallucination score of the model on $Q_{H}$ before and after training.
We explored two architectural approaches: (1) MERGED-ADAPTER, and (2) DYNAMIC-ADAPTER. In the case of MERGED-ADAPTER, the adapter was simply merged into the base model after training, effectively altering the model’s weights. Meanwhile, DYNAMIC-ADAPTER was inspired by SERAC Mitchell et al. (2022) and DAP-Adapter Ke et al. (2023); in this case, the adapter was not merged, and instead was dynamically enabled depending on the given prompt. The adapter enabling was controlled by a router-classifier model, which was trained on a mix of samples from $Q_{H}$ and $Q_{NH}$.
Table 3 and Table 4 present the full evaluation results before Self-Learning and after Self-Learning. The $Q_{H}$ dataset is the set of questions with hallucination acquired from Self-Questioning. Besides measuring the hallucination score, we also used ROUGE-Lsum Lin (2004) to measure the similarity between the model’s answer and the ground truth acquired from Knowledge Searching. Additionally, we used gpt-4o-2024-05-13 to judge the correctness of the model’s answer relative to the ground truth. The LLM-Judge was prompted to answer "yes" if the model’s answer could capture the meaning contained in the ground truth; "partly" if the model’s answer could be considered partially correct relative to the ground truth; "no" if the model’s answer was completely different from the ground truth. The responses from the LLM-Judge were then normalized: "yes" answers were converted to 1.0; "partly" answers were converted to 0.5; "no" answers, as well as any other response not following the instruction, were converted to 0.0.
Meanwhile, the Wiki dataset was acquired from Wikipedia Simple Dump 2023-12-01, serving as a test dataset to observe catastrophic forgetting. We measured the perplexity, a commonly used metric to estimate the language modeling capability of an LLM. A drastic increase of perplexity on the Wiki dataset after Self-Learning would indicate catastrophic forgetting. Finally, we also used the cleaned version of Alpaca Taori et al. (2023) as a test dataset to observe if the model would lose some knowledge related to instruction following; instructions and inputs from the dataset were used in the prompts, and then the model’s responses were compared against the expected outputs from the dataset.
Table 3: Model evaluation results before and after self-learning using MERGED-ADAPTER.
Dataset Metric BEFORE MERGED-ADAPTER
$\mathbf{Q_{H}}$ Hallucination Score 0.73 0.30
ROUGE-Lsum 0.36 0.42
LLM-Judge 0.19 0.30
Wiki Perplexity 12.02 13.11
Alpaca-Cleaned ROUGE-Lsum 0.30 0.33
LLM-Judge 0.28 0.31
Table 4: Model evaluation results before and after self-learning using DYNAMIC-ADAPTER.
Dataset Metric BEFORE DYNAMIC-ADAPTER
$\mathbf{Q_{H}}$ Hallucination Score 0.73 0.29
ROUGE-Lsum 0.36 0.42
LLM-Judge 0.19 0.30
Wiki Perplexity 12.02 12.86
Alpaca-Cleaned ROUGE-Lsum 0.30 0.33
LLM-Judge 0.28 0.32
In both approaches, despite the small size of the data and merely 3 epochs of training, we can observe greatly reduced hallucination on $Q_{H}$ after self-learning. The model has become more consistent in answering these particular questions. Furthermore, the LLM-Judge score was also increased by 11 percentage points (pp.) in both cases, from 19% to 30%. A higher number of training epochs might be needed to make the model’s answers more faithful to the ground truth and increase the LLM-Judge score further. Interestingly, the model also gained some improvement on the Alpaca dataset, with 3pp. increase using MERGED-ADAPTER and 4pp. increase using DYNAMIC-ADAPTER.
Both MERGED-ADAPTER and DYNAMIC-ADAPTER exhibited slightly increased perplexity on the Wiki dataset: 1.09 points in the former, and 0.84 points in the latter. In both cases, the increase in perplexity was very low, so catastrophic forgetting did not happen. This can be attributed to LoRA, which only changed a small number of weights in the model. MERGED-ADAPTER experienced a bigger increase in perplexity because the adapter was merged, permanently changing the affected weights. Meanwhile, with DYNAMIC-ADAPTER, the adapter was enabled only on some examples, while the original weights of the model were unaffected.
The significant reduction of hallucination on $Q_{H}$ after training proves the possibility of Self-Learning to increase the Known regions and reduce the Unknown regions of the model. Nevertheless, further investigation into long-term Self-Learning is still needed.